OpenStudy (helpmeee123):

a map is drawn using a scale of 3 centimeters = 5 meters. on the map, towab A is 7 cm from Town B. What is the actual distance between the two towns

OpenStudy (texaschic101):

3 cm 7 cm ---- = ---- 5 m x m now this is a proportion and x stands for " how many meters " so we now cross multiply 3(x) = 7(5) 3x = 35 x = 35/3 x = 11 2/3 meters so 7 cm = 11 2/3 meters (actual distance between the two towns)